OpenStudy (anonymous):

Graph the linear inequality. x + 2y = 5 I don't have a graphing calculator -_-

OpenStudy (aravindg):

ya this one is an easy task , let me guide you first divide the whole equation by 5

OpenStudy (aravindg):

@Esmy what did you get?

OpenStudy (anonymous):

x - 2y=5 x=2y+5 -3x+2y=7 2y=3x+7 3x=2y-7 2y-7 x-------- 3 2y=5= 2y-7 ------ 3

OpenStudy (anonymous):

D: ?

OpenStudy (anonymous):

but I don't have a graphing calculator so I don't know how it looks what part is shaded?

OpenStudy (amorfide):

okay rearrange this equation to get 2y= 2y=-x+5 then you will sub in x values x -2 -1 0 1 2 3 2y 7 6 5 4 3 2 so to get y you divide all x values by 2 this gives x -1 -1/2 0 1/2 1 3/2 y 7/2 3 5/2 2 3/2 1

OpenStudy (anonymous):

Can you draw how it would look :/ sry

OpenStudy (amorfide):

 something similar to this i believe

OpenStudy (amorfide):

it should go through all points but i find it hard to draw on pc

OpenStudy (anonymous):

you should have test points to know where the shaded area is.

OpenStudy (anonymous):

 Yeah it looks like that and then its shaded right?